# Recent and Upcoming Developments in Randomized Numerical Linear Algebra for Machine Learning

### v1

@@ 5.2 Statistical Inference Approaches
Bootstrap error estimation.
One application of this viewpoint is error estimation. A common short-coming of Classical RandNLA error guarantees is that they are overly pessimistic due to their worst-case nature, and therefore, they do not provide accurate predictions of the actual error incurred by the produced estimates. For instance, consider the sketched LS estimator: given a deterministic matrix $A\in\mathbb{R}^{n\times d}$ and a vector $b\in\mathbb{R}^{n}$, reduce the problem with a random sketching matrix $S\in\mathbb{R}^{m\times n}$, defining $\tilde{A}:=SA$ and $\tilde{b}:=Sb$, and use that to compute $\tilde{x}:=\operatorname*{\mathop{\mathrm{argmin}}}_{x}\|\tilde{A}x-\tilde{b}\|_{2}.$ How well does this particular estimate $\tilde{x}$ approximate the optimum $x^{*}:=\operatorname*{\mathop{\mathrm{argmin}}}_{x}\|Ax-b\|_{2}$? In other words, we would like to efficiently compute a numerical estimate $\tilde{q}(\alpha)$ such that
@@
Another application of the statistical viewpoint on RandNLA is providing a fine-grained performance analysis of different sketching or estimation methods which are indistinguishable under worst-case bounds [64, 36, 110, 52]. In particular, such analysis has been performed for the above sketched LS estimator, showing that with a fixed sketch-to-data aspect ratio $m/n$, asymptotic error for orthogonal (Hadamard or Haar) sketches is less than for i.i.d/sub-Gaussian sketches [64, 110].
@@ 5.3 Connections with Modern RMT
As an example, consider the task of LR matrix approximation. As discussed in Sec. 2.3, in this task our goal is to construct a near-optimal rank $k$ approximation of an $m\times n$ matrix $A$ by computing a $n\times k$ sketch $AS$ and then projecting the matrix $A$ onto the rank $k$ subspace defined by $AS$. Then, we can measure the quality of the output by comparing its error to the error of the best rank $k$ approximation of $A$, obtained via the Top-$k$ Truncated Singular Value Decomposition (Top-$k$ SVD)
Sketching approximation $\text{Sketching approximation factor}\quad=\quad\frac{\|A-P_{AS}A\|_{F}^{2}}{\|A-A_{k}\|_{F}^{2}},$
where $\|A-P_{AS}A\|_{F}^{2}$ is the LR projection error as in (3). Classical worst-case analysis suggests that even for sub-Gaussian sketching methods, the approximation factor for this basic approach is large and grows with $k$, as seen in Fig. 3, which has prompted computationally more expensive approaches such as power iteration [95]. However, recent work [48, 51] shows that for many data matrices, the true approximation factor is much smaller, i.e., it typically does not grow with $k$, and we can characterize it fully based on the spectral decay profile of the matrix $A$. These ideas have been further developed in other applications of RandNLA, including randomized PCA [36], Gaussian Processes [29], iterative optimization [49, 110], and in connections with the double descent phenomenon in over-parameterized ML models [52].
@@
Another important consequence of the fine-grained RMT analysis in RandNLA is a characterization of the implicit regularization phenomenon. Implicit regularization commonly refers to the regularizing effect that stochastic algorithms (such as SGD) have on the trained ML model [149, 130]; but, fundamentally, it has to do with explicit statistical problems that fast approximation algorithms implicitly solve exactly [129, 125]. Empirical results [167, 45] have shown that such a regularizing effect is caused by the stochasticity coming from the use of sketching operators; and recent theory makes this precise [52, 113].
@@
Here, again, Classical RandNLA analysis is not sensitive enough to detect the implicit regularization effect, suggesting that the correct regularization parameter should be the same as for the original task, $\tilde{\gamma}=\gamma$. However, recent fine-grained results have enabled characterizing the bias of the sketched estimators showing that we must shrink the regularization parameter $\tilde{\gamma}$ to correct for implicit regularization [45]:
@@
where $d_{\gamma}=\mathrm{tr}(A^{\scriptscriptstyle{\top}}A(A^{\scriptscriptstyle{\top}}A+\gamma I)^{-1})$ is the $\gamma$-effective dimension of $A$ and $s$ is the sketch size. Similar implicit regularization results have been shown for sketching-based distributed regression [65], sampling-based ensemble methods [113], stochastic optimization [49], and randomized rounding [63].
@@ 7 Concluding thoughts
There are many topics we did not cover: data-driven methods for learning good sketching operators [100, 101, 37]; RandNLA methods for tensor decompositions [168, 115, 77, 103]; RandNLA methods in streaming/online environments [28]; probabilistic numerics [17]; as well as topics such as Hutchinson and spectral function methods [157, 176], using randomness deep inside an algorithm, e.g., for pivot rule decisions, using these methods for the theory/practice for neural networks, e.g., NTK and Nystromformer, using these methods for quantization and low-precision computation [83, 84, 63], and using these methods to deal with graph-related and other structured matrices such as Laplacians [71, 106, 127]. In many of these cases, it is increasingly-important to identify core linear algebraic structures and build algorithmic/statistical methods around them (rather than “tacking” them on later as a “band aid” to fix problems that arise when one ignores key linear algebraic issues). This will be increasingly-important as heterogeneous hardware trends and ML training trends continue to grow. We expect the “Modern” theoretical tools developed by/for RandNLA that we have covered will be useful in these and other situations.

### v2

@@ 5.2 Statistical Inference Approaches
Bootstrap error estimation for sketching.
One application of this viewpoint is error estimation. A common short-coming of Classical RandNLA error guarantees is that they are overly pessimistic due to their worst-case nature, and therefore, they do not provide accurate predictions of the actual error incurred by the produced estimates. For instance, consider the sketched LS estimator: given a deterministic matrix $A\in\mathbb{R}^{n\times d}$ and a vector $b\in\mathbb{R}^{n}$, reduce the problem with a random sketching matrix $S\in\mathbb{R}^{s\times n}$, defining $\tilde{A}:=SA$ and $\tilde{b}:=Sb$, and use that to compute $\tilde{x}:=\operatorname*{\mathop{\mathrm{argmin}}}_{x}\|\tilde{A}x-\tilde{b}\|_{2}.$ How well does this particular estimate $\tilde{x}$ approximate the optimum $x^{*}:=\operatorname*{\mathop{\mathrm{argmin}}}_{x}\|Ax-b\|_{2}$? In other words, we would like to efficiently compute a numerical estimate $\tilde{q}(\alpha)$ such that
@@
Another application of the statistical viewpoint on RandNLA is providing a fine-grained performance analysis of different sketching or estimation methods which are indistinguishable under worst-case bounds [63, 35, 109, 51]. In particular, such analysis has been performed for the above sketched LS estimator, showing that with a fixed sketch-to-data aspect ratio $s/n$, asymptotic error for orthogonal (Hadamard or Haar) sketches is less than for i.i.d/sub-Gaussian sketches [63, 109].
@@ 5.3 Connections with Modern RMT
As an example, consider the task of LR matrix approximation. As discussed in Sec. 2.3, in this task our goal is to construct a near-optimal rank $k$ approximation of an $m\times n$ matrix $A$. A simple way of doing this involves computing an $m\times k$ sketch $AS$ and then projecting the matrix $A$ onto the rank $k$ subspace defined by $AS$. Then, we can measure the quality of the output by comparing its error to the error of the best rank $k$ approximation of $A$, obtained via the Top-$k$ Truncated Singular Value Decomposition (Top-$k$ SVD):
[…]
where $\|A-P_{AS}A\|_{F}^{2}$ is the LR projection error as in (3), and $\|A-A_{k}\|_{F}^{2}$ is the Top-$k$ SVD error. Classical worst-case analysis suggests that even for Gaussian-like sketching methods, the approximation factor for this basic approach is large and grows with $k$, as seen in Fig. 3 (black dotted line). This has been addressed by computationally more expensive approaches that involve increasing the sketch size or using block power iteration [94]. However, recent RMT-style analysis (Fig. 3, red solid line; see [47, 50]) reveals that for many data matrices, the true approximation factor is much smaller, i.e., it typically does not grow with $k$, and we can characterize it fully based on the spectral decay profile of the matrix $A$. These ideas have been further developed in other applications of RandNLA, including randomized PCA [35], Gaussian Processes [28], iterative optimization [48, 109], and in connections with the double descent phenomenon in over-parameterized ML models [51].
@@
Another important consequence of the fine-grained RMT analysis in RandNLA is a characterization of the implicit regularization phenomenon. Implicit regularization commonly refers to the regularizing effect that stochastic algorithms (such as SGD) have on the trained ML model [149, 130]; but, fundamentally, it has to do with explicit statistical problems that fast approximation algorithms implicitly solve exactly [129, 125]. Empirical results [168, 44] have shown that such a regularizing effect is caused by the stochasticity coming from the use of sketching operators; and recent theory makes this precise [51, 112, 113].
@@
Here, again, Classical RandNLA analysis is not sensitive enough to detect the implicit regularization effect, and suggests that the correct regularization parameter should be the same as for the original task, $\tilde{\gamma}=\gamma$. However, recent fine-grained results have enabled characterizing the bias of the sketched estimators showing that we must shrink the regularization parameter $\tilde{\gamma}$ to correct for implicit regularization [44]:
@@
where $d_{\gamma}=\mathrm{tr}(A^{\scriptscriptstyle{\top}}A(A^{\scriptscriptstyle{\top}}A+\gamma I)^{-1})$ is the $\gamma$-effective dimension of $A$ and $s$ is the sketch size. Similar implicit regularization results have been shown for sketching-based distributed regression [64], sampling-based ensemble methods [112], stochastic optimization [48], LR approximation [50, 113] and randomized rounding [62].
@@ 7 Concluding thoughts
There are many topics we did not cover: data-driven methods for learning good sketching operators [99, 100, 36]; RandNLA methods for tensor decompositions [169, 115, 76, 102]; RandNLA methods in streaming/online environments [27]; probabilistic numerics [16]; as well as topics such as Hutchinson and spectral function methods [158, 177], using randomness deep inside an algorithm, e.g., for pivot rule decisions, using these methods for the theory/practice of neural networks, e.g., NTK and Nystromformer, using these methods for quantization and low-precision computation [82, 83, 62], and using these methods to deal with graph-related and other structured matrices such as Laplacians [70, 105, 127]. In many of these cases, it is increasingly-important to identify core linear algebraic structures and build algorithmic/statistical methods around them (rather than “tacking” them on later as a “band aid” to fix problems that arise when one ignores key linear algebraic issues). This will be increasingly-important as heterogeneous hardware trends and ML training trends continue to grow. We expect the “Modern” theoretical tools developed by/for RandNLA that we have covered will be useful in these and other situations.
Acknowledgments
MD would like to acknowledge NSF CAREER for partial support. MWM would like to acknowledge the NSF and ONR via its BRC on RandNLA for partial support of this work.